# Magnetic properties of monolayer, multilayer, and bulk CrTe2

### v3

@@ Abstract
We investigate magnetic properties of CrTe2 within the density functional theory (DFT) approach in ferromagnetic phase and the combination of the DFT and the dynamical mean field theory (DFT+DMFT) approach in paramagnetic phase. We show that a few layer CrTe2 possesses well formed local magnetic moments. In the monolayer CrTe2 we find the most preferable antiferromagnetic exchange with the 120°$\degree$ antiferromagnetic structure. In the bilayer and trilayer systems, electronic correlations in DFT+DMFT approach yield ferromagnetic exchange interaction within each layer, but the interaction between the layers is antiferromagnetic, such that alternation of the direction of magnetization of the layers is expected. In bulk CrTe2 we find the tendency to ferromagnetic order at low temperature, but with increase of temperature antiferromagnetic correlations between the layers dominate. Determination of the critical number of layers at which the interlayer antiferromagnetic order changes to the ferromagnetic likely requires consideration of the non-local Coulomb interactions.
@@ I Introduction
Recently, chromium ditelluride CrTe2 has been suggested to exhibit ferromagnetic order even up to room temperature, making it an attractive candidate for further research in this field. Bulk CrTe2 was shown freitas_ferromagnetism_2015 to posses FM order with $T_{C}=330$ K and in-plane magnetic anisotropy. Recent years several groups obtained thin CrTe2 flakes via mechanical exfoliation purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , chemical vapor deposition meng_anomalous_2021 and molecular beam epitaxy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; liu_wafer-scale_2023 ; zheng_two-dimensional_2023 , showing stable FM order with $T_{C}\sim 200$ K. Although thick films were shown to have an in-plane magnetic anisotropy purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , the films thinner 10 monolayers show out-of-plane magnetic anisotropy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; meng_anomalous_2021 . The authors of Refs. zhang_room-temperature_2021 ; meng_anomalous_2021 ; xian_spin_2022 ; liu_wafer-scale_2023 ; wang_strain-_2024 succeeded fabricating monolayer devices. At the same time, the results of the magnetic state of monolayers remain controversial. Experimentally both, FM zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 and zigzag antiferromagnetic (AFM-zz) xian_spin_2022 states were reported.
@@
Theoretically, the density functional theory (DFT) approach was used to study various magnetic orders and exchange interactions in CrTe2. In particular, FM lv_strain-controlled_2015 ; li_tunable_2021 ; liu_structural_2022 ; yang_tunable_2021 ; yao_control_2023 , AFM gao_thickness_2021 ; AFM2 ; abuawwad_noncollinear_2022 , charge density wave otero_fumega_controlled_2020 and incommensurate abuawwad_noncollinear_2022 states were considered as possible candidates in monolayer CrTe2. It was suggested in Refs. wu_-plane_2022 ; zhu_insight_2023 ; gao_thickness_2021 ; lv_strain-controlled_2015 ; AFM2 that relatively small strain can cause the transition from AFM-zz to FM, which may explain different kinds of magnetic order observed in experiment. Furthermore, it was shown in Ref. gao_thickness_2021 that the intralayer coupling of 2 to 4 layer CrTe2 is still antiferromagnetic in DFT, similar to that of monolayer CrTe2, and changes to ferromagnetic for a larger number of layers. At the same time, the interlayer exchange interaction was also found to change from AFM to FM between 4 and 5 layers. The change of the magnetic orders was related by the authors of Ref. gao_thickness_2021 to the change of the lattice constants from monolayer to bulk. However, later study of Ref. wu_-plane_2022 has shown that at least in the bilayer system the in-plane antiferromagnetic order is stable in DFT for a wide range of lattice constants.
@@
At the same time, since Cr is an open almost half-filled $d$ shell metal, it possesses strong correlations, which may affect preferable magnetic state. Zhu et. al. zhu_insight_2023 showed that the effect of Hubbard interaction can be essential for the choice of magnetic phase in CrTe2. Also, the authors of Ref. gao_thickness_2021 emphasized that the on-site Coulomb repulsion changes the interlayer coupling to AFM even in the bulk; important role of the Coulomb repulsion for bilayer systems was also discussed in Ref. Interlayer . Previous theoretical works yao_control_2023 ; wu_-plane_2022 ; liu_structural_2022 ; li_tunable_2021 ; Interlayer used DFT$+U$ approach to account for correlation effects. However, for metallic systems this approach may be insufficient. To treat the effect of correlations, we consider in the present study the DFT+DMFT approach. We use recently proposed method of evaluation of exchange interactions in the paramagnetic phase MyJ to obtain an unbiased estimate of exchange interactions, which account for correlation effects and also consider their temperature dependence. We compare the results obtained with those from the DFT approach.
@@ II Methods and parameters
For bulk CrTe2 we choose the experimental lattice parameters freitas_ferromagnetism_2015 $a_{\rm bulk}=3.7887$, $c_{\rm bulk}=6.0955$. The momentum grid $18\times 18\times 18$ was chosen. For the monolayer CrTe2 we choose the lattice parameter of Ref. abuawwad_noncollinear_2022 $a=3.71$ and consider interlayer spacing $c=20$ to avoid overlap of the orbitals between the layers; the momentum grid $20\times 20\times 1$ was chosen. The vertical position of Te atoms was relaxed and constituted the distance $1.55$ for the bulk and $1.62$ for the monolayer system from the Cr plane. For the bilayer (trilayer) CrTe2 we choose the lattice parameter of Ref. liu_structural_2022 $a=3.76$  ($a=a_{\rm bulk}$) and consider the unit cell of vertical size $c=40$ (60 Å). The interlayer spacing was chosen equal to $c_{\rm bulk}$, Ref. sun_room_2020 . The relaxed vertical position of Te atoms constituted the distance $1.61$  ($1.56$) from the Cr plane for the outer (inner) atoms for the bilayer system and $1.58$  ($1.55$) for the trilayer system. The momentum grid was chosen $24\times 24\times 1$.
@@
To determine exchange interactions, we consider the effective Heisenberg model with the Hamiltonian $H=-(1/2)\sum_{{\bf q},rr^{\prime}}J^{rr^{\prime}}_{\bf q}{\mathbf{S}}^{r}_{\mathbf{q}}{\mathbf{S}}^{r^{\prime}}_{-{\mathbf{q}}}$, $\mathbf{S}^{r}_{\mathbf{q}}$ is the Fourier transform of static operators ${\mathbf{S}}_{ir}$ at the site $r$ of cell $i$, where the orbital-summed static spin operators ${\mathbf{S}}_{ir}=\sum_{m}{\mathbf{S}}_{irm}$ The exchange interactions in DFT were estimated by magnetic force theorem (MFT) approach korotin_calculation_2015-1 ; liechtenstein_local_1987 ; git using electron Green’s functions, calculated in Wannier basis $G_{rr^{\prime},\sigma}^{mm^{\prime}}(\mathbf{k},i\nu_{n})=\left[(i\nu_{n}+\mu)I-H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}\right]_{rm,r^{\prime}m^{\prime}}^{-1}$, where $H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}$ is the Wannier Hamiltonian, $I$ is the identity matrix, and the inversion is performed in the site- and orbital space,
@@
×$\times\left.\Delta_{r^{\prime}}\tilde{G}_{r^{\prime}r,\uparrow}({\mathbf{k}},i\nu_{n})\right]$ (1)
@@ III Results
Magnetic exchange interactions of the monolayer CrTe2 at $\beta=20$ eV-1 ($T=580$ K) are shown in Fig. 5. The maximum of exchange interaction in both, the DFT and DFT + DMFT approaches is also found at the wave vector ${\bf Q}_{K}$, which corresponds to the 120°$\degree$ antiferromagnetic order. FM order is locally stable in DFT due to the local maximum at the $\Gamma$ point with $J_{0}>0$ and unstable in DMFT, but in both cases the $\Gamma$ point is not a global maximum, which corresponds to an instability of ferromagnetism in Heisenberg model. We find that the peak of exchange interaction $J_{\mathbf{q}}$ at $\mathbf{q}=\mathbf{Q}_{K}$ in DFT approach sharply increases with a decrease of temperature (see the results of $\beta=10$ eV-1 in Fig. 5 for comparison), due to nesting of the Fermi surface pockets with the wave vector close to $\mathbf{Q}_{K}$ (see Appendix A). At the same time, the exchange interactions in the DFT+DMFT approach are quite weakly temperature-dependent because of the damping of electronic excitations. We also note that in the ferromagnetic ordered phase of DFT we obtain the magnetic moment of chromium sites $\mathfrak{m}=2.86\mu_{B}$, which corresponds to the effective spin $S_{\rm ord}=\mathfrak{m}/(g\mu_{B})=1.48$. This reflects different spin states of chromium in the ferro- and paramagnetic phases: while the chromium occupation is close in two approaches ($n_{\rm Cr}=4.67$ in DFT), the effective spin difference $\Delta S=S_{\rm eff}-S_{\rm ord}\simeq 1$ corresponds to only partial magnetization of orbital states. In particular, we find the strongest magnetization contribution $0.8\mu_{B}$ / f.u. from the $d_{3z^{2}-r^{2}}$ state and the weakest contribution $0.4\mu_{B}$ / f.u. from each of the $d_{xz,yz}$ states.
@@
DFT Energies, this work ${-1.52}$ 1.021.02{1.02}1.02 ${-0.17}$
DFT MFT, this work $-1.51$ $3.52$ $2.06$
DMFT, this work   $0.08$ $1.16$ $-1.34$
@@
all-electron KKR-GF abuawwad_noncollinear_2022 $-3.19$ $2.37$ $1.18$
@@
We also used an alternative method to obtain exchange interactions, based on the comparison of energies of various spin configurations li_tunable_2021 , see Appendix B. The comparison of the obtained exchange interactions is presented in Table 1. All DFT approaches, except Ref. li_tunable_2021 yield the same signs of nearest- and next-nearest neighbor exchange interactions, with different magnitude. In particular, the nearest neighbour exchange interaction is negative (except Ref. li_tunable_2021 ) showing the tendency to antiferromagnetism. We find the sAFM-zz configuration in DFT approach 10 meV/f.u. lower than FM. At the same time, in DMFT approach we find weak ferromagnetic nearest neighbor exchange interaction, and the tendency to antiferromagnetism originates from the next to next nearest neighbors and further interactions.
@@
We consider the interaction between the atoms of the same Cr plane (intralayer part), $J^{rr}_{\mathbf{q}}$ and the interplane interaction $J^{rr^{\prime}}_{\mathbf{q}}$ with $r\neq r^{\prime}$. From the momentum dependence of intralayer exchange interactions $J^{11}_{\mathbf{q}}=J^{22}_{\mathbf{q}}$ (see Fig. 7a) one can see that an additional sheet of the Fermi surface strongly affects exchange interactions: the intralayer interaction becomes maximal near the $\Gamma$ point of the Brillouin zone, while the height of the local maximum at the $K$ point decreases in comparison to the monolayer system. The obtained maximum of the intralayer exchange interaction can be contrasted to the results of DFT approach of the present study (shown by dashed line in Fig. 7 a) and the earlier DFT study zhu_insight_2023 , where the antiferromagnetic ground state was obtained. In case of bilayer system, both the DFT and the DFT+DMFT results for exchange interactions are weakly temperature dependent because of the absence of nesting and shift of the peak of the density of states from the Fermi level. We have also verified that the tendency to the in-plane ferromagnetic order in DFT+DMFT approach in paramagnetic phase is stable for $a>3.73$Å.
Figure 10: (Color online) Real parts of the exchange interactions $J_{ij}({\mathbf{q}})\equiv J^{ij}_{\mathbf{q}}$ in trilayer system at $\beta=18$ eV-1 (645 K) in DFT+DMFT approach. The pairs of indexes $ij$ at the exchange interactions correspond to the pairs of layers (1-bottom, 2-middle, 3-top layer)
@@ IV Conclusions
In summary, we have investigated magnetic properties of monolayer, bilayer, trilayer, and bulk CrTe2. In monolayer system in both the DFT and DFT + DMFT approaches, we find a preference for the incommensurate magnetic order with the wave vector ${\mathbf{Q}}_{K}$, which corresponds to 120°  spin alignment. This order competes with the AFM-zz magnetic order, previously suggested for monolayer system.
In the bilayer and trilayer CrTe2 in the DFT+DMFT approach, we find, in contrast to the DFT results, the tendency to the intralayer ferromagnetic order, which appears due to electronic correlations. At the same time, similarly to previous DFT$+U$ studies, we find a tendency to the interlayer AFM order. This tendency is preserved at not too low temperatures in bulk CrTe2, and changes to FM order at low temperatures.
@@ Appendix B Energy mapping of monolayer system
Figure 15: (Color online) Schematic spin configuration for (a) FM (b) AFM-ABAB (c) AFM-AABB (d) AFM-zz.

### v4

@@ Abstract
[…]
@@ I Introduction
Recently, chromium ditelluride CrTe2 has been suggested to exhibit ferromagnetic order even up to room temperature, making it an attractive candidate for further research in this field. Bulk CrTe2 was shown freitas_ferromagnetism_2015 to posses FM order with $T_{C}=330$ K and in-plane magnetic anisotropy. Recent years several groups obtained thin CrTe2 flakes via mechanical exfoliation purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , chemical vapor deposition meng_anomalous_2021 and molecular beam epitaxy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; liu_wafer-scale_2023 ; zheng_two-dimensional_2023 , showing stable FM order with $T_{C}\sim 200$ K. Although thick films were shown to have an in-plane magnetic anisotropy purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , the films thinner 10 monolayers show out-of-plane magnetic anisotropy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; meng_anomalous_2021 . The authors of Refs. zhang_room-temperature_2021 ; meng_anomalous_2021 ; xian_spin_2022 ; liu_wafer-scale_2023 ; wang_strain-_2024 succeeded fabricating monolayer devices. At the same time, the results of the magnetic state of monolayers remain controversial. Experimentally both, FM zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 and zigzag antiferromagnetic (AFM-$z$) xian_spin_2022 states were reported.
@@
Theoretically, the density functional theory (DFT) approach was used to study various magnetic orders and exchange interactions in CrTe2. In particular, FM lv_strain-controlled_2015 ; li_tunable_2021 ; liu_structural_2022 ; yang_tunable_2021 ; yao_control_2023 , AFM gao_thickness_2021 ; AFM2 ; abuawwad_noncollinear_2022 , charge density wave otero_fumega_controlled_2020 and incommensurate abuawwad_noncollinear_2022 states were considered as possible candidates in monolayer CrTe2. It was suggested in Refs. wu_-plane_2022 ; zhu_insight_2023 ; gao_thickness_2021 ; lv_strain-controlled_2015 ; AFM2 that relatively small strain can cause the transition from AFM-$z$ to FM, which may explain different kinds of magnetic order observed in experiment. Furthermore, it was shown in Ref. gao_thickness_2021 that the intralayer coupling of 2 to 4 layer CrTe2 is still antiferromagnetic in DFT, similar to that of monolayer CrTe2, and changes to ferromagnetic for a larger number of layers. At the same time, the interlayer exchange interaction was also found to change from AFM to FM between 4 and 5 layers. The change of the magnetic orders was related by the authors of Ref. gao_thickness_2021 to the change of the lattice constants from monolayer to bulk. However, later study of Ref. wu_-plane_2022 has shown that at least in the bilayer system the in-plane antiferromagnetic order is stable in DFT for a wide range of lattice constants.
@@
At the same time, since Cr is an open almost half-filled $d$ shell metal, it possesses strong correlations, which may affect preferable magnetic state. Zhu et. al. zhu_insight_2023 showed that the effect of Hubbard interaction can be essential for the choice of magnetic phase in CrTe2. Also, the authors of Ref. gao_thickness_2021 emphasized that the on-site Coulomb repulsion changes the interlayer coupling to AFM even in the bulk; important role of the Coulomb repulsion for bilayer systems was also discussed in Ref. Interlayer . Previous theoretical works yao_control_2023 ; wu_-plane_2022 ; liu_structural_2022 ; li_tunable_2021 ; Interlayer used DFT+U+U approach to account for correlation effects. However, for metallic systems this approach may be insufficient. To treat the effect of correlations, we consider in the present study the DFT+DMFT approach. We use recently proposed method of evaluation of exchange interactions in the paramagnetic phase MyJ to obtain an unbiased estimate of exchange interactions, which account for correlation effects and also consider their temperature dependence. We compare the results obtained with those from the DFT approach.
@@ II Methods and parameters
For bulk CrTe2 we choose the experimental lattice parameters freitas_ferromagnetism_2015 $a_{\rm bulk}=3.7887$ Å, $c_{\rm bulk}=6.0955$ Å. The momentum grid $18\times 18\times 18$ was chosen. For the monolayer CrTe2 we choose the lattice parameter of Ref. abuawwad_noncollinear_2022 $a=3.71$ Å and consider interlayer spacing $c=20$ Å to avoid overlap of the orbitals between the layers; the momentum grid $20\times 20\times 1$ was chosen. The vertical position of Te atoms was relaxed and constituted the distance $1.55$ Å for the bulk and $1.62$ Å for the monolayer system from the Cr plane. For the bilayer (trilayer) CrTe2 we choose the lattice parameter of Ref. liu_structural_2022 $a=3.76$ Å  ($a=a_{\rm bulk}$) and consider the unit cell of vertical size $c=40$ Å (60 Å). The interlayer spacing was chosen equal to $c_{\rm bulk}$, Ref. sun_room_2020 . The relaxed vertical position of Te atoms constituted the distance $1.61$ Å  ($1.56$ Å) from the Cr plane for the outer (inner) atoms for the bilayer system and $1.58$ Å  ($1.55$ Å) for the trilayer system. The momentum grid was chosen $24\times 24\times 1$.
@@
To determine exchange interactions, we consider the effective Heisenberg model with the Hamiltonian $H=-(1/2)\sum_{{\bf q},rr^{\prime}}J^{rr^{\prime}}_{\bf q}{\mathbf{S}}^{r}_{\mathbf{q}}{\mathbf{S}}^{r^{\prime}}_{-{\mathbf{q}}}$, $\mathbf{S}^{r}_{\mathbf{q}}$ is the Fourier transform of static operators ${\mathbf{S}}_{ir}$ at the site $r$ of cell $i$, where the orbital-summed static spin operators ${\mathbf{S}}_{ir}=\sum_{m}{\mathbf{S}}_{irm}$ The exchange interactions in DFT were estimated by magnetic force theorem (MFT) approach korotin_calculation_2015-1 ; liechtenstein_local_1987 ; git using electron Green’s functions, calculated in Wannier basis $G_{rr^{\prime},\sigma}^{mm^{\prime}}(\mathbf{k},i\nu_{n})=\left[(i\nu_{n}+\mu)I-H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}\right]_{rm,r^{\prime}m^{\prime}}^{-1}$, where $H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}$ is the Wannier Hamiltonian, II is the identity matrix, and the inversion is performed in the site- and orbital space,
@@
$\times\left.\Delta_{r^{\prime}}\tilde{G}_{r^{\prime}r,\uparrow}({\mathbf{k}},i\nu_{n})\right]$ (1)
@@ III Results
[…]
@@
DFT Energies, this work 1.52{-1.52} 1.02{1.02} 0.17{-0.17}
DFT MFT, this work 1.51-1.51 $3.52$ $2.06$
DMFT, this work $0.08$ $1.16$ 1.34-1.34
@@
all-electron KKR-GF abuawwad_noncollinear_2022 3.19-3.19 $2.37$ $1.18$
@@
We also used an alternative method to obtain exchange interactions, based on the comparison of energies of various spin configurations li_tunable_2021 , see Appendix B. The comparison of the obtained exchange interactions is presented in Table 1. All DFT approaches, except Ref. li_tunable_2021 yield the same signs of nearest- and next-nearest neighbor exchange interactions, with different magnitude. In particular, the nearest neighbour exchange interaction is negative (except Ref. li_tunable_2021 ) showing the tendency to antiferromagnetism. We find the sAFM-$z$ configuration in DFT approach 10 meV/f.u. lower than FM. At the same time, in DMFT approach we find weak ferromagnetic nearest neighbor exchange interaction, and the tendency to antiferromagnetism originates from the next to next nearest neighbors and further interactions.
@@
We consider the interaction between the atoms of the same Cr plane (intralayer part), $J^{rr}_{\mathbf{q}}$ and the interplane interaction $J^{rr^{\prime}}_{\mathbf{q}}$ with $r\neq r^{\prime}$. From the momentum dependence of intralayer exchange interactions $J^{11}_{\mathbf{q}}=J^{22}_{\mathbf{q}}$ (see Fig. 7a) one can see that an additional sheet of the Fermi surface strongly affects exchange interactions: the intralayer interaction becomes maximal near the $\Gamma$ point of the Brillouin zone, while the height of the local maximum at the $K$ point decreases in comparison to the monolayer system. The obtained maximum of the intralayer exchange interaction can be contrasted to the results of DFT approach of the present study (shown by dashed line in Fig. 7 a) and the earlier DFT study zhu_insight_2023 , where the antiferromagnetic ground state was obtained. In case of bilayer system, both the DFT and the DFT+DMFT results for exchange interactions are weakly temperature dependent because of the absence of nesting and shift of the peak of the density of states from the Fermi level. We have also verified that the tendency to the in-plane ferromagnetic order in DFT+DMFT approach in paramagnetic phase is stable for a>3.73a>3.73Å.
Figure 10: (Color online) Real parts of the exchange interactions $J_{ij}({\mathbf{q}})\equiv J^{ij}_{\mathbf{q}}$ in trilayer system at $\beta=18$ eV-1 (645 K) in DFT+DMFT approach. The pairs of indexes ijij at the exchange interactions correspond to the pairs of layers (1-bottom, 2-middle, 3-top layer)
@@ IV Conclusions
In summary, we have investigated magnetic properties of monolayer, bilayer, trilayer, and bulk CrTe2. In monolayer system in both the DFT and DFT + DMFT approaches, we find a preference for the incommensurate magnetic order with the wave vector ${\mathbf{Q}}_{K}$, which corresponds to 120°  spin alignment. This order competes with the AFM-$z$ magnetic order, previously suggested for monolayer system.
In the bilayer and trilayer CrTe2 in the DFT+DMFT approach, we find, in contrast to the DFT results, the tendency to the intralayer ferromagnetic order, which appears due to electronic correlations. At the same time, similarly to previous DFT+U+U studies, we find a tendency to the interlayer AFM order. This tendency is preserved at not too low temperatures in bulk CrTe2, and changes to FM order at low temperatures.
@@ Appendix B Energy mapping of monolayer system
Figure 15: (Color online) Schematic spin configuration for (a) FM (b) AFM-ABAB (c) AFM-AABB (d) AFM-$z$.